Consider an ideal diatomic gas that performs a reversible thermodynamic cycle that consists of a succession of the following transformations.

A → B: isobar expansion from volume VA to the volume VB = 2 · VA;

B → C: isochoric transformation;

C → D: isobar expansion up to the volume VD = 3 · VA;

D → A: isothermal compression that closes the cycle.

Calculate the efficiency of this reversible cycle


1
Expert's answer
2021-12-20T10:30:56-0500

"\eta=1-\frac{Q_1}{Q_2}=1-\frac{p\Delta V_2}{p\Delta V_1}=1-\frac{\Delta V_2}{\Delta V_1}=\frac 23=67\%."